# Nature of Topological Phase Transition of Kitaev Quantum Spin Liquids

### v1

@@ Abstract
The spin-$1/2$ honeycomb Kitaev model [1] [Fig. 1(a)] has been at the forefront of research into quantum spin liquids (QSLs) [2, 3, 4] since it is exactly solvable after fractionalizing the spin operators into a set of Majorana fermions [1, 5, 6]. Some of these correspond to local bond excitations which are linked to $Z_{2}$ fluxes through the plaquettes of the honeycomb lattice. Since the fluxes are conserved, the Kitaev model can be diagonalized for each flux configuration, resulting in a non-interacting Hamiltonian for the remaining Majorana fermion species. In the ground state, zero-flux sector, this results in a Dirac dispersion identical to that of electrons in graphene.
Anisotropy of the Kitaev couplings can drive a topological phase transition from a gapless to a gapped $Z_{2}$ Kitaev QSL [1]. In the regime of large anisotropy, the latter can be mapped to the toric code model which exhibits anyonic excitations and plays an important role in the context of quantum computation and quantum error correction [7]. Approaching the topological phase transition from the gapless QSL side, the Dirac points of the gapless Majorana bands move along the edge of the Brillouin zone [Fig. 1(b)] and eventually merge, forming a semi Dirac point with a quadratic and a linear band touching direction. For larger anisotropies the spectrum becomes gapped. This behavior is very similar to the topological phase transition proposed to occur for electrons in strained honeycomb lattices [8, 9, 10] and was observed experimentally in black phosphorus [11, 12].
[…]
@@
Figure 2: Mean-field phase diagram as a function of the anisotropy $(K_{z}-K)/K$ and the Ising exchange $J/K$. The evolution of the Majorana fermion spectrum across the topological phase transition between the gapless and gapped quantum spin liquid phases is shown in the insets.
Even for the pure Kitaev model, $J=0$, this seems to be an interacting problem. However, in this case the $\hat{\eta}$ Majorana fermions only live on isolated $z$ bonds and the bond operators $\hat{B}_{z}(\mathbf{r})=i\hat{\eta}_{A}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r})$, which have eigenvalues ±$\pm 1$, commute with the Hamiltonian, $[\hat{B}_{z}(\mathbf{r}),\hat{\cal H}]=0$. In the absence of flux excitations we can replace all operators $\hat{B}_{z}(\mathbf{r})$ with the negative eigenvalue. This results in a non-interacting Hamiltonian for the $\hat{\psi}$ Majorana fermions with energy dispersion $\epsilon_{\psi,\pm}(\mathbf{k})=\pm|K_{z}+K(e^{i\mathbf{k}\cdot\mathbf{a}_{1}}+e^{i\mathbf{k}\cdot\mathbf{a}_{2}})|$. For $K_{z}/K<2$ we obtain gapless excitations with a pair of Dirac points. These merge at $K_{z}/K=2$ into a semi-Dirac point at $\mathbf{K}_{s}=(0,\frac{2\pi}{3})$. For $K_{z}/K>2$ the spectrum is gapped.
For non-zero $J$ the $\hat{\eta}$ Majorana fermions acquire dynamics and $[\hat{B}_{z}(\mathbf{r}),\hat{\cal H}]\neq 0$. In this case the model is no longer exactly solvable. An approximate phase diagram of the Kitaev-Ising model can be obtained using mean-field theory [27], where the bond expectation values $A_{\gamma}=\langle i\hat{\psi}_{A}(\mathbf{r})\hat{\psi}_{B}(\mathbf{r}+\bm{\delta}_{\gamma})\rangle$ and $B_{\gamma}=\langle i\hat{\eta}_{A}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r}+\bm{\delta}_{\gamma})\rangle$ ($\bm{\delta}_{x}=\mathbf{a}_{1}$, $\bm{\delta}_{y}=\mathbf{a}_{2}$, $\bm{\delta}_{z}=\bm{0}$), as well as the staggered magnetization $m=\langle i\hat{\psi}_{A}(\mathbf{r})\hat{\eta}_{A}(\mathbf{r})\rangle=-\langle i\hat{\psi}_{B}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r})\rangle$ are determined self-consistently. This results in the phase diagram shown in Fig. 2.
As expected, a relatively small Ising exchange $J$ leads to a first-order transition to an antiferromagnetic state. Importantly, a continuous topological phase transition between a gapless and a gapped Kitaev QSL still occurs for sufficiently small $J$. The insets of Fig. 2 show the evolution of the mean-field dispersion across this transition. While the gapless $\hat{\psi}$ Majorana modes behave in the same way as for the pure anisotropic Kitaev model, a key difference is that the gapped $\hat{\eta}$ modes become dispersive.
In order to understand the nature of the topological quantum phase transition, it is essential to include fluctuations beyond mean-field theory, arising from the interaction vertex. We recast the problem using a Grassmann path integral with action
@@
where $\tau$ denotes imaginary time, $\mathbf{k}$ the two-dimensional momentum, $k_{0}$ frequency, and $k=(k_{0},\mathbf{k})$. The complex functions $\xi_{\mathbf{k}}=\sum_{\gamma}a_{\gamma}e^{i\mathbf{k}\bm{\delta}_{\gamma}}$ and $\lambda_{\mathbf{k}}=\sum_{\gamma}b_{\gamma}e^{i\mathbf{k}\bm{\delta}_{\gamma}}$ are linked to the mean-field dispersions, $\epsilon_{\psi,\pm}(\mathbf{k})=\pm|\xi_{\mathbf{k}}|$ and $\epsilon_{\eta,\pm}(\mathbf{k})=\pm|\lambda_{\mathbf{k}}|$, respectively. We have written the interactions as $g_{\gamma}$, for brevity. Because of symmetry $g_{x}=g_{y}$, $a_{x}=a_{y}$ and $b_{x}=b_{y}$. Note that $b_{z}/b_{x,y}>2$ since the $\eta$ Majorana fermion bands are gapped.
@@
$S_{Y}[\phi,\bar{\bm{\psi}},\bm{\psi}]=\frac{g}{\sqrt{N}}\int_{k,q}\phi(q)\bar{\bm{\psi}}(k)\bm{\sigma}_{y}\bm{\psi}(k+q),$ (11)
where $\bm{\sigma}_{y}$ denotes a Pauli matrix in sublattice space and the Ising fluctuation field is given by $\frac{g}{\sqrt{N}}\phi(q)=\Omega_{z}(q)-\Omega_{x}(q)-\Omega_{y}(q)$. Note that we generalized to $N$ flavours of semi-Dirac fermions and scaled the coupling accordingly. Expanding the quadratic part $S_{0}[\bar{\bm{\psi}},\bm{\psi}]$ of the action (8) around $\mathbf{K}_{s}$ we obtain
$S_{0}[\bar{\bm{\psi}},\bm{\psi}]=\int_{k}\bar{\bm{\psi}}_{k}\left[-ik_{0}+k_{L}\bm{\sigma}_{x}+\left(k_{Q}^{2}+\Delta\right)\bm{\sigma}_{y}\right]\bm{\psi}_{k},$ (12)
@@
with $G_{\phi}^{-1}(q)=\Pi(q)$ as starting point for subsequent perturbative RG calculation [28]. Using the correct infrared (IR) scaling form of the propagator, the fluctuation corrections under RG are independent of the choice of the ultraviolet (UV) cut-off scheme and therefore universal [29]. The bosonic self energy $\Pi(q)=g^{2}/N\int_{k}\text{Tr}\left[G_{\bm{\psi}}(k)\bm{\sigma}_{y}G_{\bm{\psi}}(k+q)\bm{\sigma}_{y}\right]$ is obtained by calculating the fermion polarization bubble digram [Fig. 4(a)] over the full range of frequencies and momenta where the non-analyticity arrises from the IR contribution ($k\to 0$). Unfortunately, for the case of semi-Dirac fermions this integral cannot be computed analytically. Following the procedure in Ref. [29], we obtain
@@
×$\times\left[\frac{(p+1)^{4}+p^{2}(p+1)^{2}+(1-t)u}{(p+1)^{4}t+p^{4}(1-t)+t(1+t)u}-2\right].\quad$
@@
The field theory $S_{0}[\bar{\bm{\psi}},\bm{\psi}]+S_{0}[\phi]+S_{Y}[\phi,\bar{\bm{\psi}},\bm{\psi}]$ for the topological phase transition between the gapless and gapped Kitaev QSL states is very similar to the GNY theory that describes the quantum criticality of semi-Dirac fermions in 2+1 dimensions due to spontaneous symmetry breaking [29, 30, 31, 32, 33]. A key difference, however, is that for the symmetry-breaking transitions the Yukawa coupling is through the $\bm{\sigma}_{z}$ channel, which upon condensation of the order parameter results in the opening of a conventional mass gap in the fermion spectrum. The different form of the Yukawa coupling (11) through $\bm{\sigma}_{y}$ changes the form of the IR propagator $G_{\phi}(q)$ and of perturbative RG diagrams, resulting in distinct critical behavior.
@@
The fermionic self energy correction $\bm{\Sigma}(k)d\ell=-g^{2}/N\int_{q}^{>}G_{\phi}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(k+q)\bm{\sigma}_{y}$, which corresponds to the diagram in Fig. 4(b), is of the same form as the original kernel in $S_{0}[\bar{\bm{\psi}},\bm{\psi}]$,
@@
where the coefficients are evaluated as [34]
@@
From the diagram shown in Fig. 4(c) be obtain the correction $\Omega\bm{\sigma}_{y}d\ell=(g^{2}/N)\int_{q}^{>}G_{\phi}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(q)\bm{\sigma}_{y}$ to the Yukawa coupling matrix, where the shell integral gives $\Omega=\Sigma_{\Delta}$.
@@
$\frac{1}{2}-\frac{0.0674}{N}$ $1-\frac{0.5823}{N}$ $\frac{0.2185}{N}$ $-\frac{1.0298}{N}$
@@
To summarize, we have derived the effective field theory for the topological quantum phase transition between the gapless and gapped Kitaev QSL phases. For the pure, exactly solvable Kitaev model the problem reduces to a free-fermion field theory. Away from the exactly solvable point, the field theory is of the GNY type and describes the coupling between an Ising fluctuation field to the gapless semi-Dirac Majorana fermion modes. We determined the critical exponents from an RG analysis and demonstrated that the universality of the topological phase transition is different to that describing symmetry-breaking phase transitions of semi-Dirac fermions.
The exponent $\nu_{\Delta}$ is linked to the opening of the energy gap in the Majorana fermion spectrum, $\Delta\sim(\delta-\delta_{c})^{\nu_{\Delta}}$, as well as to the separation of the Dirac points on the gapless QSL side, $(\delta k)\sim(\delta_{c}-\delta)^{\nu_{\Delta}/2}$. It could in principle be determined experimentally by measuring the evolution of the magnetic excitation continuum across the topological phase transition. However, it remains a challenge to realize the topological phase transition in experiment since uniaxial strain would not only affect the anisotropy of the Kitaev couplings but also distort the lattice, resulting in an increase of other magnetic exchange couplings.

### v2

@@ Abstract
The spin-$1/2$ honeycomb Kitaev model [1] [Fig. 1(a)] has been at the forefront of research into quantum spin liquids (QSLs) [2, 3, 4, 5, 6] since it is exactly solvable after fractionalizing the spin operators into a set of Majorana fermions [1, 7, 8]. Some of these correspond to local bond excitations which are linked to $Z_{2}$ fluxes through the plaquettes of the honeycomb lattice. Since the fluxes are conserved, the Kitaev model can be diagonalized for each flux configuration, resulting in a non-interacting Hamiltonian for the remaining Majorana fermion species. In the ground state, zero-flux sector, this results in a Dirac dispersion identical to that of electrons in graphene.
Anisotropy of the Kitaev couplings can drive a topological phase transition from a gapless to a gapped $Z_{2}$ Kitaev QSL [1]. In the regime of large anisotropy, the latter can be mapped to the toric code model which exhibits anyonic excitations and plays an important role in the context of quantum computation and quantum error correction [9]. Approaching the topological phase transition from the gapless QSL side, the Dirac points of the gapless Majorana bands move along the edge of the Brillouin zone [Fig. 1(b)] and eventually merge, forming a semi Dirac point with a quadratic and a linear band touching direction. For larger anisotropies the spectrum becomes gapped. This anisotropy-driven topological phase transition is not characterized by Chern numbers of the bands in the gapped state, unlike the topological phase transitions driven by magnetic fields [10, 11]. It is instead similar to the topological phase transition of real electrons in strained honeycomb lattices [12, 13, 14], which was observed experimentally in black phosphorus [15, 16].
At first glance, the bond-directional exchange of the Kitaev model seems artificial, but it was realized that because of strong spin-orbital mixing [17, 18], the Kitaev model can be approximately realized in layered honeycomb iridates [19, 20, 21, 22, 23, 24, 25] and the halide $\alpha$-RuCl3 [26, 27, 28]. Although in these materials the additional magnetic interactions are still slightly too large, leading to magnetic ordering at low temperatures, the experimental realization of a Kitaev QSL is certainly within reach.
@@
Figure 2: Mean-field phase diagram as a function of the anisotropy $(K_{z}-K)/K$ and the Ising exchange $J/K$. The evolution of the Majorana fermion spectrum across the topological phase transition between the gapless and gapped quantum spin liquid phases is shown in the insets. Adapted from Ref. [31].
Even for the pure Kitaev model, $J=0$, this seems to be an interacting problem. However, in this case the $\hat{\eta}$ Majorana fermions only live on isolated $z$ bonds and the bond operators $\hat{B}_{z}(\mathbf{r})=i\hat{\eta}_{A}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r})$, which have eigenvalues ±$\pm 1$, commute with the Hamiltonian, $[\hat{B}_{z}(\mathbf{r}),\hat{\cal H}]=0$. In the absence of flux excitations, corresponding to the ground-state sector [32], we can replace all operators $\hat{B}_{z}(\mathbf{r})$ with the negative eigenvalue. This results in a non-interacting Hamiltonian for the $\hat{\psi}$ Majorana fermions with energy dispersion $\epsilon_{\psi,\pm}(\mathbf{k})=\pm|K_{z}+K(e^{i\mathbf{k}\cdot\mathbf{a}_{1}}+e^{i\mathbf{k}\cdot\mathbf{a}_{2}})|$. For $K_{z}/K<2$ we obtain gapless excitations with a pair of Dirac points. These merge at $K_{z}/K=2$ into a semi-Dirac point at $\mathbf{K}_{s}=(0,\frac{2\pi}{3})$. For $K_{z}/K>2$ the spectrum is gapped.
For non-zero $J$ the $\hat{\eta}$ Majorana fermions acquire dynamics and $[\hat{B}_{z}(\mathbf{r}),\hat{\cal H}]\neq 0$. In this case the model is no longer exactly solvable. An approximate phase diagram of the Kitaev-Ising model can be obtained using mean-field theory [31], where the bond expectation values $A_{\gamma}=\langle i\hat{\psi}_{A}(\mathbf{r})\hat{\psi}_{B}(\mathbf{r}+\bm{\delta}_{\gamma})\rangle$ and $B_{\gamma}=\langle i\hat{\eta}_{A}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r}+\bm{\delta}_{\gamma})\rangle$ ($\bm{\delta}_{x}=\mathbf{a}_{1}$, $\bm{\delta}_{y}=\mathbf{a}_{2}$, $\bm{\delta}_{z}=\bm{0}$), as well as the staggered magnetization $m=\langle i\hat{\psi}_{A}(\mathbf{r})\hat{\eta}_{A}(\mathbf{r})\rangle=-\langle i\hat{\psi}_{B}(\mathbf{r})\hat{\eta}_{B}(\mathbf{r})\rangle$ are determined self-consistently. This results in the phase diagram shown in Fig. 2, which is adapted from Ref. [31].
As expected, a small Ising exchange $J$ leads to a first-order transition to a magnetically ordered state [31, 33, 30]. Importantly, a continuous topological phase transition between a gapless and a gapped Kitaev QSL still occurs for sufficiently small $J$. The insets of Fig. 2 show the evolution of the mean-field dispersion across this transition. While the gapless $\hat{\psi}$ Majorana modes behave in the same way as for the pure anisotropic Kitaev model, a key difference is that the gapped $\hat{\eta}$ modes become dispersive.
In order to understand the nature of the topological quantum phase transition, it is essential to include fluctuations beyond mean-field theory, arising from the interaction vertex. We recast the problem using a Grassmann path integral with action [34]
@@
where $\tau$ denotes imaginary time, $\mathbf{k}$ the two-dimensional momentum, $k_{0}$ frequency, and $k=(k_{0},\mathbf{k})$. The fields $\bm{\psi}_{k}$ and $\bm{\eta}_{k}$ are two component spinors in sublattice space and we have written $\bm{\psi}^{\dagger}_{k}=(\psi_{A}(-k),\psi_{B}(-k))$, for brevity. The complex functions $\xi_{\mathbf{k}}=\sum_{\gamma}a_{\gamma}e^{i\mathbf{k}\bm{\delta}_{\gamma}}$ and $\lambda_{\mathbf{k}}=\sum_{\gamma}b_{\gamma}e^{i\mathbf{k}\bm{\delta}_{\gamma}}$ are linked to the mean-field dispersions, $\epsilon_{\psi,\pm}(\mathbf{k})=\pm|\xi_{\mathbf{k}}|$ and $\epsilon_{\eta,\pm}(\mathbf{k})=\pm|\lambda_{\mathbf{k}}|$, respectively. We have written the interactions as $g_{\gamma}$, for brevity. Because of symmetry $g_{x}=g_{y}$, $a_{x}=a_{y}$ and $b_{x}=b_{y}$. Note that $b_{z}/b_{x,y}>2$ since the $\eta$ Majorana fermion bands are gapped.
@@
$S_{Y}[\phi,\bm{\psi}]=\frac{g}{\sqrt{N}}\int_{k,q}\phi(q)\bm{\psi^{\dagger}}(k)\bm{\sigma}_{y}\bm{\psi}(k+q),$ (11)
where $\bm{\sigma}_{y}$ denotes a Pauli matrix in sublattice space and the Ising fluctuation field is given by $\frac{g}{\sqrt{N}}\phi(q)=\Omega_{z}(q)-\Omega_{x}(q)-\Omega_{y}(q)$. Note that we generalized from two ($A$, $B$) to an even number of $N$ Majorana fermion flavors and scaled the coupling accordingly. Expanding the quadratic part $S_{0}[\bm{\psi}]$ of the action (8) around $\mathbf{K}_{s}$ we obtain
$S_{0}[\bm{\psi}]=\int_{k}\bm{\psi}^{\dagger}_{k}\left[-ik_{0}+k_{L}\bm{\sigma}_{x}+\left(k_{Q}^{2}+\Delta\right)\bm{\sigma}_{y}\right]\bm{\psi}_{k},$ (12)
@@
with $G_{\phi}^{-1}(q)=\Pi(q)$ as starting point for subsequent perturbative RG calculation [35]. Using the correct infrared (IR) scaling form of the propagator, the fluctuation corrections under RG are independent of the choice of the ultraviolet (UV) cut-off scheme and therefore universal [36]. The bosonic self energy $\Pi(q)=g^{2}/N\int_{k}\text{Tr}\left[G_{\bm{\psi}}(k)\bm{\sigma}_{y}G_{\bm{\psi}}(k+q)\bm{\sigma}_{y}\right]$ is obtained by calculating the fermion polarization bubble digram [Fig. 4(a)] over the full range of frequencies and momenta where the non-analyticity arrises from the IR contribution ($k\to 0$). Unfortunately, for the case of semi-Dirac fermions this integral cannot be computed analytically. As shown in the Supplemental Material [37], we obtain
@@
×$\times\left[\frac{(p+1)^{4}+p^{2}(p+1)^{2}+(1-t)u}{(p+1)^{4}t+p^{4}(1-t)+t(1-t)u}-2\right].\quad$
Eq. (14) smoothly connects the asymptotic forms $\Pi(q)\sim|q_{Q}|$ for $q_{0}=q_{L}=0$ and $\Pi(q)\sim(q_{0}^{2}+q_{L}^{2})^{1/4}$ for $q_{Q}=0$.
@@
The field theory $S_{0}[\bm{\psi}]+S_{0}[\phi]+S_{Y}[\phi,\bm{\psi}]$ for the topological phase transition between the gapless and gapped Kitaev QSL states is very similar to the GNY theory that describes the quantum criticality of semi-Dirac fermions in 2+1 dimensions due to spontaneous symmetry breaking [36, 38, 39, 40, 41]. A key difference, however, is that for the symmetry-breaking transitions the Yukawa coupling is through the $\bm{\sigma}_{z}$ channel, which upon condensation of the order parameter results in the opening of a conventional mass gap in the fermion spectrum. The different form of the Yukawa coupling (11) through $\bm{\sigma}_{y}$ changes the form of the IR propagator $G_{\phi}(q)$ and of perturbative RG diagrams, resulting in distinct critical behavior.
@@
The fermionic self energy correction $\bm{\Sigma}(k)d\ell=-g^{2}/N\int_{q}^{>}G_{\phi}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(k+q)\bm{\sigma}_{y}$, which corresponds to the diagram in Fig. 4(b), is of the same form as the original kernel in $S_{0}[\bm{\psi}]$,
@@
where the coefficients are evaluated as (see Supplemental Material [37])
@@
From the diagram shown in Fig. 4(c) be obtain the correction $\Omega\bm{\sigma}_{y}d\ell=(g^{2}/N)\int_{q}^{>}G_{\phi}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(q)\bm{\sigma}_{y}G_{\bm{\psi}}(q)\bm{\sigma}_{y}$ to the Yukawa coupling matrix, where the shell integral gives $\Omega=\Sigma_{\Delta}$ (see Supplemental Material [37]).
@@
$\frac{1}{2}-\frac{0.0337}{N}$ $1-\frac{0.2911}{N}$ $\frac{0.1092}{N}$ $-\frac{0.5149}{N}$
@@
To summarize, we have derived the effective field theory for the topological quantum phase transition between the gapless and gapped Kitaev QSL phases. For the pure, exactly solvable Kitaev model the problem reduces to a free-fermion field theory. Away from the exactly solvable point, the field theory is of the GNY type and describes the coupling between an Ising fluctuation field to the gapless semi-Dirac Majorana fermion modes. We determined the critical exponents from an RG analysis and demonstrated that the universality of the topological phase transition is different to that describing symmetry-breaking phase transitions of semi-Dirac electrons.
It is also important to compare with the nature of the topolological phase transition of semi-Dirac electrons subject to long-range Coulomb repulsions [35]. In this case the Hubbard-Stratonovich decoupling results in a GNY theory with a bosonic fluctuation field that couples to the local density of electrons. Note that such density operators do not exist for a single Majorana fermion mode. As we have demonstrated, the fluctuation fields in Kitaev QSLs instead couple to bond operators of pairs of Majorana fermions, resulting in an off-diagonal Yukawa coupling matrix proportional to $\sigma_{y}$. Because of this crucial difference the IR boson propagator acquires a different form and the critical exponents are different.
Let us briefly discuss potential links with experiments. Tuning across the topological phase transition one would expect to see crossovers in the specific heat at temperatures much smaller than the energy gap of flux excitations. While the separated Dirac points in the gapless Kitaev QSL result in a quadratic temperature dependence of the specific heat, $C\sim T^{2}$, the distinct density of states of semi-Dirac fermions would give rise to a $C\sim T^{3/2}$ dependence above the critical point. Finally, in the gapped Kitaev QSL phase the specific heat is exponentially suppressed, $C\sim e^{-T/\Delta}$. As suggested in the literature [42], a quantum critical fan-shape temperature dependence across the zero-temperature topological quantum phase transition could be uncovered through measurements of thermal Hall conductivity.
Unlike for electronic Dirac semimetals, the bandstructure of the emergent Majorana fermions of Kitaev QSLs cannot be directly probed experimentally. However, the magnetic excitation continua measured by inelastic neutron scattering are linked to fermionic particle-hole excitations, making it possible, in principle, to extract the exponent $\nu_{\Delta}$. The topological phase transition results in the opening of an energy gap in the Majorana-fermion spectrum, $\Delta\sim(\delta-\delta_{c})^{\nu_{\Delta}}$, corresponding to an energy gap $2\Delta$ of the magnetic excitation continuum. Approaching the topological phase transition from the gapless QSL side the separation of the Dirac points vanishes as $\delta k=\sqrt{-\Delta}\sim(\delta_{c}-\delta)^{\nu_{\Delta}/2}$. For this momentum transfer particle-hole excitations with zero energy are possible, resulting in a gap closing of the magnetic excitation continuum at $\delta k$. One might also anticipate signatures of the topological phase transition in the quantum Fisher information [43] which is a witness for quantum entanglement and can be calculated from the measured dynamical susceptibility [44].